OpenStudy (anonymous):

Can anyone explain why 3^3/5 x 3^ 8/5 = 3^11/5 instead of 9^11/5

OpenStudy (amistre64):

bases arent "numbers" in that sense

OpenStudy (anonymous):

because 8+3 = 11

OpenStudy (anonymous):

Because they are powers. When multiplying with powers you add the exponents.

OpenStudy (amistre64):

they are symbols that represent another number

OpenStudy (anonymous):

for the same reason that \[10^2\times 10^3=10^5\] and not \[10^2\times 10^3=100^5\]

OpenStudy (amistre64):

3^2 = 9 4^3 = 64

OpenStudy (amistre64):

the base is not a "number", its a representation of a number

OpenStudy (anonymous):

hmm

OpenStudy (amistre64):

3^2 * 3^3 means: (3*3)*(3*3*3) = (3*3*3*3*3) = 3^5

OpenStudy (anonymous):

Yeah I get that

OpenStudy (amistre64):

as you can see; 3^5 doesnt mean 9^5 does it?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

You know what. Just think of it this way, when you multiply powers, you don't ACTUALLY multiply all the numbers that are present. You just need to add the exponents. When you're dividing, just do subtraction with the exponents.

OpenStudy (anonymous):

So if I had 3^2/3 x 4^2/3 what would my answer be?

OpenStudy (amistre64):

well, handwaving never amounts to understanding does it.

OpenStudy (amistre64):

youd have no answer in that form; since they represent 2 unlike terms

OpenStudy (anonymous):

You have to have the same base number.

OpenStudy (amistre64):

3 is not the same term as 4

OpenStudy (amistre64):

you could still come to an answer; just not a shortcut answer :)

OpenStudy (anonymous):

so why can't you multiply the bases if the bases have exponents

OpenStudy (amistre64):

we just established why.

OpenStudy (amistre64):

3*3*3*3*3 not= 9*9*9*9*9 ...

OpenStudy (amistre64):

the rules havent changed any fromthen til now :)

OpenStudy (amistre64):

spose i develop a system of math that say RW$/ can add to *)♫i and produce any prime number i want

OpenStudy (amistre64):

all that means is that I have established a rule that has a certain consequence and rests soley on the definition of the symbols

OpenStudy (amistre64):

bases and exponents are just like that; there are only certain rules that will make them useful

OpenStudy (anonymous):

I think I see what you are saying, so if I had 3^2 x 4^3 I could not solve that until the exponents are gone

OpenStudy (amistre64):

correct; or unless you could convert them in some fashion to like bases

OpenStudy (anonymous):

oh okay I think I get it now! Thank you guys!

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (anonymous):

What grade are you in?

OpenStudy (anonymous):

I'd rather not say lol...

OpenStudy (anonymous):

goodbye

OpenStudy (anonymous):

That's okay. hahaa.